• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name R. Michaels

    Log entry time 00:08:38 on July14,2005

    Entry number 147997

    This entry is a followup to: 147990

    keyword=T0 adjusted

    I adjusted the L1A timing (T0) on each arm so that the T0's of
    the VDCs line up with past history, and this fixed several problems except
    still no S2m TDCs on R-arm. I'll need to check with a
    scope and understand & finalize the T0 setup.

    Run 2037 is in progress with the new timing.